myininaya (myininaya):

Sally can paint a house in 4 hours. John can paint the same house in 6 hours. How long will it take them to paint it together?

OpenStudy (anonymous):

12/5 hours

OpenStudy (anonymous):

1/4+1/6=1/x

OpenStudy (anonymous):

5\12 hrs

OpenStudy (anonymous):

form the eqn, n solve for x

myininaya (myininaya):

fruitless is right! gj lol

myininaya (myininaya):

i did x(1/4+1/6)=1 which is basically what you have up there uzma so i should give you a medal

OpenStudy (anonymous):

3+2/12=1/x 1/x=5/12 x=12/5

myininaya (myininaya):

vg

myininaya (myininaya):

lol

OpenStudy (anonymous):

:)

myininaya (myininaya):

should i give you a more challenging question lol

OpenStudy (anonymous):

challenging like this one;)??

myininaya (myininaya):

lol

OpenStudy (anonymous):

now i dont exactly remember how many times this house had been painted:P

OpenStudy (anonymous):

\[\frac{x}{4}+\frac{x}{6}=1\] denominator is 24 \[6x+4x=24\] \[10x=24\] \[x=24: 10\] \[x=2.4 hours\] or : 2 hours 24 minute